# Joint Discriminative-Generative Modeling via Dual Adversarial Training

### v1

@@ A.4.2 Data augmentation details
Figure 3: Training curves under different data augmentations during stage 2 joint training.
A.5 Additional analyses
A.5.1 Individual contributions of generative loss and decoupled augmentation
The results in Table˜7 demonstrate the impacts of the AT-based generative loss and decoupled augmentation. Introducing the generative loss component to the standard AT baseline significantly reduces the FID from 33.04 to 15.35, while robust accuracy remains comparable. The subsequent application of a decoupled augmentation strategy yields a further reduction in FID to 9.07.
Table 7: Effect of generative loss and augmentation on CIFAR-10.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$
Standard AT 92.34 75.73 33.04
DAT (uniform aug) 92.68 75.93 15.35
DAT (decoupled aug) 91.86 75.66 9.07
RATIO [Augustin et al., 2020] 92.23 76.25 21.96
A.5.2 Effect of OOD dataset size
Table 8: Impact of OOD dataset size on ImageNet performance for DAT ResNet-50 ($T=15$).
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$ IS $\uparrow$
Standard AT 62.83 34.44 15.97 274.90
DAT 1K w/o aug 57.50 33.80 8.00 320.64
DAT 1K 57.56 34.22 6.94 324.23
DAT 10K 57.82 34.70 6.84 320.78
DAT 100K 58.19 34.88 6.70 322.10
DAT 300K 57.88 34.84 6.64 339.55
A.5.3 Generative-discriminative trade-off via loss weighting
Our training objective, $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$, is a composite of a discriminative and a generative loss. This structure naturally raises the question of whether it is possible to trade off between these two capabilities by adjusting the relative weight of each component. To investigate this possibility, we perform experiments on CIFAR-10 with three distinct weighting configurations:
  • Standard loss (equal weighting): $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$
  • Emphasize generative modeling: $\mathcal{L}(\theta)=0.6\cdot\mathcal{L}_{\text{AT-CE}}(\theta)+1.4\cdot\mathcal{L}_{\text{BCE}}(\theta)$
  • Emphasize classification: $\mathcal{L}(\theta)=1.4\cdot\mathcal{L}_{\text{AT-CE}}(\theta)+0.6\cdot\mathcal{L}_{\text{BCE}}(\theta)$
The results in Table 9 confirm that the balance between generative and discriminative performance can be tuned by adjusting the loss term weights. Emphasizing the generative component improves FID at the cost of slightly reduced classification performance, while emphasizing classification achieves the opposite effect. However, we note that our standard, unweighted loss corresponds to the natural factorization of the joint log-likelihood in the original JEM formulation: $\log p_{\theta}(x,y)=\log p_{\theta}(y|x)+\log p_{\theta}(x)$. This suggests that equal weighting is a principled default that performs well without requiring additional hyperparameter tuning.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$
Standard loss 91.88 75.73 9.09
Emphasize generative modeling 91.16 75.11 8.77
Emphasize classification 92.52 75.97 10.02
A.6 Generative performance evaluation
@@
where $T$ is the number of PGD steps and $\eta$ is the corresponding step size (see Table˜10).
@@
The FID results for both conditional and unconditional generation across all datasets are presented in Table˜11. We find conditional generation consistently outperforms unconditional generation across all the datasets.
@@
Model Dataset PGD steps ($T$) Step size OOD data source DAT CIFAR-10 ($T=40$) 33 0.2 80M Tiny Images CIFAR-10 ($T=50$) 35 0.2 80M Tiny Images CIFAR-100 ($T=45$) 32 0.2 80M Tiny Images CIFAR-100 ($T=50$) 33 0.2 80M Tiny Images ImageNet (ResNet-50, $T=15$) 13 8.0 Open Images ImageNet (ResNet-50, $T=30$) 16 8.0 Open Images ImageNet (WRN-50-4, $T=30$) 17 8.0 Open Images ImageNet (WRN-50-4, $T=65$) 20 8.0 Open Images RATIO CIFAR-10 31 0.2 80M Tiny Images CIFAR-100 14 0.2 80M Tiny Images Standard AT CIFAR-10 22 0.2 80M Tiny Images CIFAR-100 15 0.2 80M Tiny Images ImageNet (ResNet-50) 13 8.0 Open Images ImageNet (WRN-50-4) 11 8.0 Open Images
@@
CIFAR-10 CIFAR-100 ImageNet Conditional generation $9.07$ $10.70$ $6.64$ Unconditional generation $20.57$ $13.56$ $18.67$
A.7 Out-of-distribution detection
We evaluate both standard out-of-distribution (OOD) detection performance and worst-case OOD detection under adversarial perturbations. For standard OOD detection, we measure the AUROC scores between in-distribution test samples and unmodified OOD samples. For worst-case detection, we evaluate against adversarially perturbed OOD samples specifically optimized to maximize the OOD detection function output. Results are computed using all the in-distribution test samples and 1024 out-distribution samples. For generating adversarial OOD samples, we use $L_{2}$-based perturbation limit of 1.0 for CIFAR-10/100 and 3.0 for ImageNet.
@@ A.7 Out-of-distribution detection
Table˜12 presents a comparison of the above two OOD detection functions. The results reveal complementary strengths: the energy-based function ($E_{\theta}(x)$) achieves near-perfect AUROC scores on uniform noise detection, while the maximum confidence function ($\max_{y}p_{\theta}(y|x)$) demonstrates superior performance on natural image OOD datasets. Based on these findings, we adopt the maximum confidence score for subsequent comparisons with other methods.
Table˜13 presents comparative results across different baselines. Notably, our DAT model achieves OOD detection performance comparable to standard AT on natural image datasets (CIFAR-100, SVHN), despite incorporating an additional OOD dataset during training. This observation suggests that our generative training component primarily enhances generation quality rather than improving OOD detection capabilities beyond those provided by standard adversarial training.
@@
A.8 Calibration
@@
A.9 Counterfactual generation
Refer to captionRefer to caption
A.10 Gneration results
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT ($T=50$).
(c) Uncurated conditional samples of RATIO.
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT ($T=50$).
(c) Uncurated conditional samples of RATIO.
Figure 6: CIFAR-100 conditional generation results.
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT (WRN-50-4 $T=65$).
(c) Uncurated conditional samples of standard AT (WRN-50-4).
Figure 7: ImageNet class-conditional generation results for the first 10 classes: tench, goldfish, great white shark, tiger shark, hammerhead, electric ray, stingray, cock, hen, ostrich (images are in $224\times 224$ resolution).
Figure 8: Selected ImageNet conditional generation results for class 88 (macaw), 107 (jellyfish), 130 (flamingo), 145 (king penguin), 248 (husky), 258 (Samoyed), 291 (lion), 511 (convertible), and 980 (volcano). Results are generated with DAT (WRN-50-4, $T=65$) at $224\times 224$ resolution.
A.11 Variability of DAT performance across datasets
@@
Table 16: Mean and standard deviation of DAT performance across datasets, computed over three independent runs with different random seeds.
@@
Dataset Acc% $\uparrow$ Robust Acc% $\uparrow$ IS $\uparrow$ FID $\downarrow$ CIFAR-10 ($T=40$) 91.86 $\pm$ 0.03 75.66 $\pm$ 0.01 9.96 $\pm$ 0.02 9.07 $\pm$ 0.03 CIFAR-100 ($T=45$) 65.55 $\pm$ 0.62 45.97 $\pm$ 0.49 10.83 $\pm$ 0.11 10.70 $\pm$ 0.22 ImageNet (ResNet-50, $T=15$) 57.91 $\pm$ 0.09 34.87 $\pm$ 0.09 334.11 $\pm$ 8.96 6.60 $\pm$ 0.08
A.12 Training curves
(a) CIFAR-10 ($T=40$)
(b) CIFAR-100 ($T=45$)
(c) ImageNet (ResNet-50, $T=15$)

### v3

@@ B.1.2 Data augmentation
Figure 4: CIFAR-10 training curves under different data augmentations during stage 2 joint training.
B.2 Computational efficiency
Training overhead relative to standard AT. We estimate computational cost as follows: a forward pass costs 1 unit, an input-gradient backward (for PGD) costs 1 unit, and a full backward (for parameter updates) costs 2 units. This gives 2 units per PGD step and 3 units per training update. Under this model, Stage 1 (standard AT) costs $(2K+3)B$ units per iteration: $2KB$ for the $K$-step PGD attack plus $3B$ for the training update. Stage 2 adds $T$ generative PGD steps ($2TB$ units) and processes three sample types per iteration (adversarial in-distribution, adversarial OOD, and clean in-distribution), tripling the training update cost to $9B$ units, for a total of $(2K+2T+9)B$ units per iteration. With $E_{1}$ and $E_{2}$ epochs for each stage, the total overhead relative to standard AT is:
$\text{Overhead}=1+\frac{E_{2}}{E_{1}}\cdot\frac{2K+2T+9}{2K+3}$ (29)
[…]
Table 7: Computational cost and performance metrics of DAT models. Training overhead computed relative to standard AT using Equation˜29. Stage 2 training times are effective durations excluding evaluation. Baseline training times converted from V100-days; throughput measured on a single MI300 accelerator.
[…]
B.3 Training dynamics
We present the stage 2 training curves to illustrate the training dynamics.
(a) CIFAR-10 ($T=40$)
(b) CIFAR-100 ($T=45$)
(c) ImageNet $\times256$ (ResNet-50, $T=15$)
(d) ImageNet $\times256$ (ConvNeXt-L, $T=110$)
B.4 Evaluation methodology
@@
where $T$ is the number of PGD steps and $\eta$ is the corresponding step size (see Table˜8).
@@
The FID results for both conditional and unconditional generation across all datasets are presented in Table˜9. We find conditional generation consistently outperforms unconditional generation across all the datasets.
@@
Model Dataset PGD steps ($T$) Step size OOD data source DAT CIFAR-10 ($T=40$) 33 0.2 80M Tiny Images CIFAR-10 ($T=50$) 35 0.2 80M Tiny Images CIFAR-100 ($T=45$) 32 0.2 80M Tiny Images CIFAR-100 ($T=50$) 33 0.2 80M Tiny Images ImageNet (ResNet-50, $T=15$) 14 8.0 Open Images ImageNet (ResNet-50, $T=30$) 14 8.0 Open Images ImageNet (WRN-50-4, $T=30$) 17 8.0 Open Images ImageNet (WRN-50-4, $T=65$) 19 8.0 Open Images ImageNet (ConvNeXt-L-CvSt, $T=110$) 36 8.0 Open Images RATIO CIFAR-10 31 0.2 80M Tiny Images CIFAR-100 14 0.2 80M Tiny Images Standard AT CIFAR-10 22 0.2 80M Tiny Images CIFAR-100 15 0.2 80M Tiny Images ImageNet (ResNet-50) 13 8.0 Open Images ImageNet (WRN-50-4) 11 8.0 Open Images ImageNet (ConvNeXt-L-CvSt) 0 8.0 Open Images
@@
CIFAR-10 CIFAR-100 ImageNet $\times224$ (ResNet50, $T=15$) Conditional generation $9.07$ $10.70$ 6.64 Unconditional generation $20.57$ $13.56$ 18.67
Appendix C Results & Analyses
C.1 Qualitative samples
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT ($T=50$).
(c) Uncurated conditional samples of RATIO.
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT ($T=50$).
(c) Uncurated conditional samples of RATIO.
Figure 7: CIFAR-100 conditional generation results.
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT (WRN-50-4 $T=65$).
(c) Uncurated conditional samples of standard AT (WRN-50-4).
Figure 8: ImageNet class-conditional generation results for the first 10 classes: tench, goldfish, great white shark, tiger shark, hammerhead, electric ray, stingray, cock, hen, ostrich (images are in $\times256$ resolution).
Figure 9: Selected ImageNet conditional generation results for class 88 (macaw), 107 (jellyfish), 130 (flamingo), 145 (king penguin), 248 (husky), 258 (Samoyed), 291 (lion), 511 (convertible), and 980 (volcano). Results are generated with DAT ConvNeXt-L-CvSt at $\times256$
C.2 Visual counterfactuals
C.3 OOD detection
We evaluate both standard out-of-distribution (OOD) detection performance and worst-case OOD detection under adversarial perturbations using models trained with $\times224$ generation. For standard OOD detection, we measure the AUROC scores between in-distribution test samples and unmodified OOD samples. For worst-case detection, we evaluate against adversarially perturbed OOD samples specifically optimized to maximize the OOD detection function output. Results are computed using all the in-distribution test samples and 1024 out-distribution samples. For generating adversarial OOD samples, we use $L_{2}$-based perturbation limit of 1.0 for CIFAR-10/100 and 3.0 for ImageNet.
@@ C.3 OOD detection
Table˜10 presents a comparison of the above two OOD detection functions. The results reveal complementary strengths: the energy-based function ($E_{\theta}(x)$) achieves near-perfect AUROC scores on uniform noise detection, while the maximum confidence function ($\max_{y}p_{\theta}(y|x)$) demonstrates superior performance on natural image OOD datasets. Based on these findings, we adopt the maximum confidence score for subsequent comparisons with other methods.
Table˜11 presents comparative results across different baselines. Notably, our DAT model achieves OOD detection performance comparable to standard AT on natural image datasets (CIFAR-100, SVHN), despite incorporating an additional OOD dataset during training. This observation suggests that our generative training component primarily enhances generation quality rather than improving OOD detection capabilities beyond those provided by standard adversarial training.
@@
C.4 Calibration
We assess model calibration using reliability diagrams on models trained with $\times224$ generation. Figures 12, 13, and 14 show calibration diagrams for CIFAR-10, CIFAR-100, and ImageNet respectively.
@@ C.4 Calibration
C.5 Corruption robustness
To evaluate whether the joint training objective maintains robustness under distribution shift, we assess our CIFAR-10 models on CIFAR-10-C [Hendrycks and Dietterich, 2019], a benchmark testing robustness to 15 common corruption types across 5 severity levels.
As shown in Table˜14, standard AT achieves a mean corruption error (mCE) of 19.63%, while DAT with $T=40$ achieves mCE of 19.84%, and DAT with $T=50$ achieves mCE of 21.84%. While there is a trade-off between generative quality and corruption robustness as $T$ increases, DAT overall maintains strong corruption robustness comparable to standard AT across all corruption types.
Table 14: Robustness evaluation on CIFAR-10. We report clean accuracy, robust accuracy, and corruption robustness on CIFAR-10-C. For corruptions, we report error rate (%) averaged across 5 severity levels. mCE denotes mean corruption error across all 15 corruption types.
Metric Standard AT DAT ($T=40$) DAT ($T=50$)
Noise corruptions (error %)
   Gaussian noise 21.30 20.52 20.63
   Shot noise 17.14 16.36 16.86
   Impulse noise 23.80 23.54 24.72
Blur corruptions (error %)
   Defocus blur 11.08 11.69 13.13
   Glass blur 15.38 14.74 17.06
   Motion blur 13.65 14.27 15.84
   Zoom blur 11.72 12.39 13.89
Weather corruptions (error %)
   Snow 12.19 12.19 13.62
   Frost 12.44 12.13 13.60
   Fog 24.75 26.67 29.45
   Brightness 8.26 8.60 9.61
Digital corruptions (error %)
   Contrast 33.31 32.89 36.24
   Elastic transform 12.34 13.06 14.80
   Pixelate 9.51 9.65 11.08
   JPEG compression 9.56 9.73 11.14
Mean corruption error (mCE) $\downarrow$ 19.63 19.84 21.84
Clean Acc (%) $\uparrow$ 92.43 91.86 90.72
Robust Acc (%) $\uparrow$ 75.73 75.66 74.65
FID $\downarrow$ 28.41 9.07 7.19
C.6 L-infinity generalization
All our models in Tables˜1 and 2 are trained with $L_{2}$-based adversarial attacks for both the discriminative and generative components. To demonstrate that our approach generalizes to $L_{\infty}$ adversarial training, we train on ImageNet $\times256$ with ConvNeXt-L using $L_{\infty}$ attacks: for discriminative training, we use 2-step PGD with $\epsilon=4/255$ and step size 2/255; for generative training, we use step size 3/255 with $T=110$ maximum steps. All other hyperparameters (optimizer, learning rate, weight decay, EMA, batch size) match our $L_{2}$ configuration exactly (see Table˜3). FID is evaluated on samples generated using $L_{\infty}$-based PGD with 36 steps and step size 0.03 (approx. 8/255).
Table˜15 presents the results. Compared to Standard AT, our $L_{\infty}$-trained DAT model trades modest reductions in clean accuracy (76.58% vs 78.25%) and robust accuracy (57.94% vs 59.40%) for significantly superior generation quality (FID 4.11 vs 44.46, IS 320.7 vs 27.32). This demonstrates that DAT successfully achieves joint discriminative-generative modeling under $L_{\infty}$ training, maintaining competitive robustness while enabling high-quality generation.
[…]
Table 15: $L_{\infty}$ and $L_{2}$ results on ImageNet $\times256$ with ConvNeXt-L. Standard AT results are based on $L_{\infty}$-trained ($\epsilon=4/255$) checkpoint from Singh et al. [2023]. Both DAT models initialize from the same $L_{\infty}$-trained Stage 1 checkpoint [Singh et al., 2023] (no $L_{2}$-trained checkpoint is available) and differ only in Stage 2 training: one uses $L_{\infty}$ perturbations, the other uses $L_{2}$ perturbations. All models are evaluated under both $L_{\infty}$ ($\epsilon=4/255$) and $L_{2}$ ($\epsilon=3.0$) adversarial attacks using AutoAttack [Croce and Hein, 2020].
Method Training Norm Clean Acc% $\uparrow$ Robust Acc ($L_{\infty}$ 4/255) $\uparrow$ Robust Acc ($L_{2}$ 3.0) $\uparrow$ FID $\downarrow$ IS $\uparrow$
Standard AT $L_{\infty}$ 78.25 59.40 33.38 44.46 27.32
DAT ($T=110$) $L_{\infty}$ 76.58 57.94 33.40 4.11 320.7
DAT ($T=110$) $L_{2}$ 75.73 51.90 56.40 3.29 310.2
C.7 Ablation studies
In this section, we dissect the contributions of individual model components, analyze the data efficiency of our generative learning mechanism, and investigate the inherent trade-off between discriminative and generative performance.
C.7.1 Component ablation
The results in Table˜16 demonstrate the impacts of the AT-based generative loss and decoupled augmentation. Introducing the generative loss component to the standard AT baseline significantly reduces the FID from 33.04 to 15.35, while robust accuracy remains comparable. The subsequent application of a decoupled augmentation strategy yields a further reduction in FID to 9.07.
Table 16: Effect of generative loss and augmentation on CIFAR-10.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$
Standard AT 92.34 75.73 33.04
DAT (uniform aug) 92.68 75.93 15.35
DAT (decoupled aug) 91.86 75.66 9.07
RATIO [Augustin et al., 2020] 92.23 76.25 21.96
C.7.2 OOD data efficiency
Table 17: Impact of OOD dataset size on ImageNet performance for DAT ResNet-50 ($T=15$) with $\times224$ generation.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$ IS $\uparrow$
Standard AT 62.83 34.44 15.97 274.90
DAT 1K w/o aug 57.50 33.80 8.00 320.64
DAT 1K 57.56 34.22 6.94 324.23
DAT 10K 57.82 34.70 6.84 320.78
DAT 100K 58.19 34.88 6.70 322.10
DAT 300K 57.88 34.84 6.64 339.55
C.7.3 Discriminative-generative trade-off
Our empirical analyses across multiple dimensions—varying PGD steps $T$ (Section˜4.3), adjusting loss weights (see below), and modifying augmentation strategies (Section˜B.1.2)—reveal an inherent trade-off between generative and discriminative performance. The augmentation ablation provides insight into a potential underlying mechanism: when all other settings remain fixed and only the augmentation strategy for the generative component varies, using no augmentation achieves better FID but worse robust accuracy, while random cropping maintains similar FID while improving robustness. This sensitivity to the generative pipeline’s data distribution suggests that model representations adapt to the data used for generative modeling. The substantially lower FID of DAT compared to standard AT is consistent with the learned representations exhibiting stronger alignment with the original data distribution $p_{\text{data}}$ than with the augmented distribution used for discriminative training. These observations suggest a tension between two objectives: the generative objective encourages representations aligned with $p_{\text{data}}$ for high-fidelity generation, while the discriminative objective benefits from alignment with the augmented distribution for robustness. While this fundamental tension cannot be fully resolved within the current framework, the trade-off can be tuned through mechanisms such as the number of PGD iterations $T$ (Equation˜8) and loss weighting to suit different application requirements.
Control via loss weighting. To explicitly investigate the controllability of this trade-off, we perform experiments on CIFAR-10 with three distinct weighting configurations for the composite objective $\mathcal{L}(\theta)=\lambda_{1}\mathcal{L}_{\text{AT-CE}}(\theta)+\lambda_{2}\mathcal{L}_{\text{BCE}}(\theta)$: (1) standard loss with $\lambda_{1}=\lambda_{2}=1.0$; (2) emphasize generative with $\lambda_{1}=0.6,\lambda_{2}=1.4$; and (3) emphasize classification with $\lambda_{1}=1.4,\lambda_{2}=0.6$. The results in Table 18 confirm that the balance between generative and discriminative performance can be tuned by adjusting the loss term weights. Emphasizing the generative component improves FID at the cost of slightly reduced classification performance, while emphasizing classification achieves the opposite effect. However, we note that our standard, unweighted loss corresponds to the natural factorization of the joint log-likelihood in the original JEM formulation: $\log p_{\theta}(x,y)=\log p_{\theta}(y|x)+\log p_{\theta}(x)$. This suggests that equal weighting is a principled default that performs well without requiring additional hyperparameter tuning.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$
Standard loss 91.88 75.73 9.09
Emphasize generative modeling 91.16 75.11 8.77
Emphasize classification 92.52 75.97 10.02
C.8 Reproducibility and stability
@@
Table 19: Mean and standard deviation of DAT performance across datasets, computed over five independent runs with different random seeds. We observed zero divergences across all runs.
@@
Dataset Acc% $\uparrow$ Robust Acc% $\uparrow$ IS $\uparrow$ FID $\downarrow$ CIFAR-10 (WRN-34-10, $T=40$) 91.92 $\pm$ 0.09 75.75 $\pm$ 0.07 9.92 $\pm$ 0.05 9.12 $\pm$ 0.05 CIFAR-100 (WRN-34-10, $T=45$, LR=0.009) 65.76 $\pm$ 0.75 45.94 $\pm$ 0.48 10.99 $\pm$ 0.29 10.73 $\pm$ 0.25 ImageNet $\times256$ (ResNet-50, $T=15$) 61.31 $\pm$ 0.16 39.96 $\pm$ 0.41 322.65 $\pm$ 2.28 6.87 $\pm$ 0.05